# Quantitative surgery and total mean curvature

### v1

@@ 5. Surgery bending in arbitrary backgrounds
$\mathrm{scal}(\Sigma_{\varrho_{1}},u_{s}^{\frac{4}{n-2}}\cdot e^{\varrho_{1}}g_{\varrho_{1}})={}$ $e^{-\varrho_{1}}u_{s}^{-\frac{n+2}{n-2}}\left(\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})u_{s}-\frac{4(n-1)}{n-2}(\Delta_{g}u_{s})\right)$
$={}$ $e^{-\varrho_{1}}u_{s}^{-\frac{n+2}{n-2}}\Bigl((1-s)\underbrace{\left(\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})u-\frac{4(n-1)}{n-2}(\Delta_{g}u)\right)}_{=u^{(n+2)/(n-2)}\mathrm{scal}\left(\Sigma_{\varrho_{1}},g_{\varrho_{1},\mu}\right)}$
$\qquad\qquad\qquad\qquad\qquad+2s\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})\Bigr).$
@@
$\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1},\mu})\geq\mathrm{scal}(M,g)-\varrho_{1}^{1/4}\geq-|\mathrm{scal}(M,g)|-\varrho_{1}^{1/4}.$
@@
Furthermore, the proof of Theorem 5.1 reveals that $\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})\geq-|\mathrm{scal}(M,g)|-\varrho_{1}^{1/4}$ as well, see (5.6), (5.7) and (5.8). Combining this with $u_{s}\geq 1$ which gives $u_{s}^{-(n+2)/(n-2)}\geq-1$, we obtain
@@
$\mathrm{scal}(\Sigma_{\varrho_{1}},u_{s}^{\frac{4}{n-2}}\cdot e^{\varrho_{1}}g_{\varrho_{1}})\ \geq{}$ $-u_{s}^{-\frac{n+2}{n-2}}e^{-\varrho_{1}}(|\mathrm{scal}(M,g)|+\varrho_{1}^{1/4})$
$\geq{}$ $-e^{-\varrho_{1}}(|\mathrm{scal}(M,g)|+\varrho_{1}^{1/4})$
which is bounded independently of $\mu$. ∎
6. Proof of Theorem 1.2
Let $\sigma\in\mathbb{R}$, $(M,g_{M})$ and $(\Omega,g_{\Omega})$ be as in Theorem 1.2. Using the arguments from the beginning of Section 2, we can assume without loss of generality that $\Omega$ is orientable and admits a spin structure for $\dim M\geq 7$. Furthermore, we can assume that $S^{n}$ can be obtained from $M$ by a sequence of surgeries of codimension at least $3$. Let $\ell$ be the minimal number of surgeries required. The proof is by induction on $\ell$, where the case $\ell=0$ has been treated in [25].
By a small deformation of $g_{\Omega}$ near $\partial\Omega\subset\Omega$, leaving the induced metric on $\partial\Omega$ unchanged and decreasing the total mean curvature and scalar curvature by an arbitrarily small amount, we can assume (replacing $\sigma$ with $\sigma-1$) that $H_{g_{\Omega}}>0$ along $\partial\Omega$. In this situation the proof of Section 2 no longer applies, since the width of the attached handles becomes smaller when $H_{g_{\Omega}}$ approaches 0. Instead, we use the following argument in the surgery step.
@@
Let $\widetilde{M}$ be obtained from $M$ by a surgery in codimension $k\geq 3$ such that $S^{n}$ can be obtained from $\widetilde{M}$ by $\ell$ surgeries of codimension at least $3$. Put d:=nkd:=n-k.
Let $\psi\colon S^{d}\times D^{k}\hookrightarrow M$ be the surgery datum for $\widetilde{M}$. Then, there exists a bordism $T\colon M\leadsto\widetilde{M}$ consisting of a single $\ell$-handle $D^{d+1}\times S^{k}$, which is glued onto $M$ along $\psi$.
Let $\mu>0$ and consider the metric $g_{\mu}$ from Theorem 5.1 on $M\times[0,1]$. This extended metric can then be further extended to a metric $\widetilde{g}_{\mu}$ on the handle $M\times[0,1]\cup_{M\times\{1\}}T$ in an arbitrary fashion. By Theorem 5.1 and Proposition 5.6, there exist constants $0<\varrho_{0}<\varrho_{1}<\frac{1}{10}$, a family of submanifolds $\Sigma_{\varrho}\subset\widetilde{\Omega}$ for $\varrho\in(0,\varrho_{1}]$ and a family of diffeomorphisms $\Phi_{\varrho}\colon\Sigma_{\varrho_{1}}\to\Sigma_{\varrho_{0}}$ such that
  1. (i)
    $\Sigma_{\varrho}$ is diffeomorphic to $\widetilde{M}$ and divides $M\times[0,1]\cup T$ into two manifolds $\widetilde{T}$ and $\widetilde{U}$, where $\widetilde{T}$ is diffeomorphic to $M\times[0,1]\cup T$.
  2. (ii)
    $\mathrm{scal}(\Sigma_{\varrho})>\mathrm{scal}(M)-\varrho^{1/4}$ for all $\varrho$.
  3. (iii)
    $H_{g_{\mu}}|_{\Sigma_{\varrho_{0}}}>0$.
  4. (iv)
    $\Phi_{\varrho_{1}}=\operatorname{\mathrm{id}}_{\Sigma_{\varrho_{1}}}$.
  5. (v)
    $\varrho\mapsto e^{\varrho}\Phi_{\varrho}^{*}(g_{\mu}|_{\Sigma_{\varrho}})$ is strictly increasing.
Note that apriori the surfaces $\Sigma_{\varrho}$ constructed in Section 5 have a boundary, however, they can be capped off in a standard fashion while preserving all geometric bounds.
@@
We also note that $\varrho_{1}$ is independent of $\mu$. By Proposition 5.6, there exists an increasing path $g_{t}$ from $e^{\varrho_{1}}\widetilde{g}|_{\Sigma_{\varrho_{1}}}$ to $2^{4/(n-2)}{g^{\prime}}|_{\Sigma_{\varrho_{1}}}$, where $g^{\prime}$ is some fixed extension of the metric $g\cup g|_{M}+\mathrm{d}t^{2}$. Furthermore, the metric $g^{\prime}$ is independent of $\mu$ and the scalar curvature of every metric of this path can be uniformly bounded independent of $\mu$.
Let us abbreviate $\Sigma\coloneqq\Sigma_{\varrho_{1}}$. We will now apply the criterion provided by Proposition 3.2. Let
$g_{\Sigma}$ $\coloneqq 2^{4/(n-2)}{g^{\prime}}|_{\Sigma}$
$s$ $\coloneqq\min\{\min(\mathrm{scal}(g_{t})),\mathrm{scal}(g)-\varrho_{1}^{1/4}\}.$
Then, by the above, $\Phi_{\varrho}^{\ast}(\widetilde{g}_{\mu}|_{\Sigma_{\varrho}})\in\mathscr{R}_{g_{\Sigma},s}(\Sigma_{\varrho_{1}})$ for every $\varrho\leq\varrho_{1}$ (see Section 3 for the definition of $\mathscr{R}_{g_{\Sigma},s}$).
First, we note that by induction hypothesis, the theorem holds for the metric $g_{\Sigma}$ on $\Sigma$. If $(\Omega,g_{\Omega})$ is a fill-in as in Theorem 1.2, then $\Omega\cup\widetilde{T}$ is a fill-in of $\Sigma$. Since we assumed $H_{g_{\Omega}}>0$ and $M$ is compact, there exists a positive lower bound $0<\mu<H_{g_{\Omega}}$. Moreover, by performing a smoothing procedure similar to the one in the proof of Theorem 2.1 we can extend the metric $\widetilde{g}_{\mu}$ to this fill-in. Since the Riemannian manifolds $M$ and $\Sigma_{\varrho_{0}}$ only differ on a small neighbourhood, we may assume that
$\int_{M}H_{g_{\Omega}}{\rm dvol}_{g_{M}}\leq 2\int_{\Sigma_{\varrho_{0}}}H_{\widetilde{g}_{\mu}}{\rm dvol}_{\widetilde{g}_{\mu}}|_{\Sigma_{\varrho_{0}}}.$
The extended metric $\widetilde{g}$ on $\widetilde{\Omega}\coloneqq\Omega\cup T$ constitutes a fill-in of $\Sigma_{\varrho_{0}}$ with positive mean curvature. Since $(\Sigma_{\varrho_{0}},\widetilde{g}|_{\Sigma_{\varrho_{0}}})$ is isometric to $(\Sigma_{\varrho_{1}},\Phi_{\varrho}^{\ast}(\widetilde{g}|_{\Sigma_{\varrho}})$ and $\Phi_{\varrho}^{\ast}(\widetilde{g}|_{\Sigma_{\varrho}})\in\mathscr{R}_{g_{\Sigma},s}$, Proposition 3.2 implies that there exists a constant $\Lambda=\Lambda(\Sigma,g_{\Sigma})$ such that
$\int_{\Sigma_{\varrho_{0}}}H_{\widetilde{g}_{\mu}}{\rm dvol}_{\widetilde{g}_{\mu}}|_{\Sigma_{\varrho_{0}}}\leq\Lambda.$
Since $\Lambda$ is independent of $\mu$ and hence independent of the fill-in, this finishes the proof.

### v2

@@ 6. Surgery bending in arbitrary backgrounds
$\mathrm{scal}(\Sigma_{\varrho_{1}},e^{\varrho_{1}}u_{s}^{\frac{4}{n-2}}\cdot g_{\varrho_{1}})={}$ $e^{-\varrho_{1}}u_{s}^{-\frac{n+2}{n-2}}\left(\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})u_{s}-\frac{4(n-1)}{n-2}(\Delta_{g_{\varrho_{1}}}u_{s})\right)$
$={}$ $e^{-\varrho_{1}}u_{s}^{-\frac{n+2}{n-2}}\Bigl((1-s)\underbrace{\left(\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})u-\frac{4(n-1)}{n-2}(\Delta_{g_{\varrho_{1}}}u)\right)}_{=u^{(n+2)/(n-2)}\mathrm{scal}\left(\Sigma_{\varrho_{1}},g_{\varrho_{1},\kappa}\right)}$
$\qquad\qquad\qquad\qquad\qquad+2s\cdot\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})\Bigr).$
@@
$\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1},\kappa})\geq\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}-\varrho_{1}^{1/4}\geq-|\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}|-\varrho_{1}^{1/4}.$
@@
Furthermore, the proof of Theorem 6.1 reveals that the same inequalities hold for $\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})$ as well, see (6.7), (6.10) and (6.12). Thus, we obtain
$\mathrm{scal}(\Sigma_{\varrho_{1}},$ $e^{\varrho_{1}}\cdot u_{s}^{\frac{4}{n-2}}\cdot g_{\varrho_{1}})$
@@
$\geq{}$ $-e^{-\varrho_{1}}u_{s}^{-\frac{n+2}{n-2}}\underbrace{\left((1-s)\cdot u^{\frac{n+2}{n-2}}+s\cdot 2\right)}_{\leq((1-s)\cdot u+s\cdot 2)^{{(n+2)}/{(n-2)}}}(|\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}|+\varrho_{1}^{1/4})$
$\geq{}$ $-(|\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}|+\varrho_{1}^{1/4})\qed$
Remark 6.8.
If the scalar curvature of $g_{M}$ is positive, we can arrange for $\mathrm{scal}(g_{s})$ to be positive in Proposition 6.7: If $\varrho_{1}$ is chosen such that $\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}-\varrho_{1}^{1/4}>0$, then $\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1}})$ and $\mathrm{scal}(\Sigma_{\varrho_{1}},g_{\varrho_{1},\kappa})$ can be bounded by $\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}-\varrho_{1}^{1/4}>0$. Since $(1-s)\cdot u^{\frac{n+2}{n-2}}+s\cdot 2\geq u_{s}$, we obtain
$\mathrm{scal}(\Sigma_{\varrho_{1}},g_{s})\geq e^{-\varrho_{1}}u_{s}^{-\frac{4}{n-2}}(\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}-\varrho^{1/4})>0.$
Remark 6.9.
Theorem 6.1 has an analogue for initial data sets $(M,g_{M},k)$, if we replace the scalar curvature by the dominant energy scalar $\mu-|J|$, where
$\mu=\frac{1}{2}(\mathrm{scal}_{g_{\Omega}}+\operatorname{tr}(k)^{2}-|k|^{2}),\qquad J=\operatorname{div}(k-\operatorname{tr}(k)g).$
That is, if we extend $k$ trivially to $M\times[0,1]$ and multiply it by a cutoff-function supported on $0,2a_{\varrho}]$, the dominant energy scalar of the hypersurface $\Sigma_{\varrho}$ only decreases by $\varrho^{1/4}$. If one goes through the proof, one encounters additional error terms for $\tfrac{1}{2}(\operatorname{tr}_{g_{M}}(k)^{2}-|k|^{2})$, $J$ and $\operatorname{tr}_{g_{\varrho}}(k)$, which have exactly the same type as the error terms already appearing for $\mathrm{scal}$ and $H$, see Remark 6.3.
7. Non-spin fill-ins with non-negative mean curvature
In this section we will prove Theorem (B) by induction on the number of surgeries required to transform $M$ into a sphere. The following is the main result needed for the induction step.
Proposition 7.1.
@@ 7. Non-spin fill-ins with non-negative mean curvature
Let $(M,g_{M})$ be a closed oriented Riemannian manifold. Let $\widetilde{M}$ be a closed oriented manifold such that $\widetilde{M}$ is obtained from $M$ by performing an oriented surgery in codimension at least $3$. Then there exists
  • a Riemannian metric $g_{\widetilde{M}}$ on $\widetilde{M}$,
  • a smooth bordism $T_{1}$ from $M$ to $\widetilde{M}$,
  • a constant $C\leq 0$,
with the following property: Let $(\Omega,g_{\Omega})$ be a fill-in of $(M,g_{M})$ with $\mathrm{scal}_{g_{\Omega}}\geq\sigma$ and $H_{g_{\Omega}}\geq 0$. Then there exists a smooth Riemannian metric $g_{\Omega_{1}}$ on
$\Omega_{1}=\Omega\bigcup_{\partial\Omega\approx M}T_{1}$
such that the following holds:
  • $g_{\Omega_{1}}=g_{\Omega}$ outside an arbitrarily small neighborhood of $T_{1}\subset\Omega_{1}$,
  • $g_{\Omega_{1}}$ induces the metric $g_{\widetilde{M}}$ on $\partial\Omega_{1}=\widetilde{M}\times\{1\}$,
  • $H_{g_{\Omega_{1}}}>0$ along $\widetilde{M}\times\{1\}$,
  • $\mathrm{scal}_{g_{\Omega_{1}}}\geq\min\{\sigma,C\}-1$,
  • $\int_{\partial\Omega}H_{g_{\Omega}}{\rm dvol}_{g_{\partial\Omega}}<2\int_{\partial\Omega_{1}}H_{g_{\Omega_{1}}}{\rm dvol}_{g_{\partial\Omega_{1}}}$.
Proof.
Let $\psi\colon S^{n-k}\times D^{k}\hookrightarrow M$, $k\geq 3$, be the surgery datum for passing from $M$ to $\widetilde{M}$. Let $\mathcal{S}:=\psi(S^{n-k}\times 0)\subset M$ be the corresponding surgery sphere. The trace of $\psi$ is defined as
$T=T(\psi)\coloneqq(M\times[0,1])\cup_{\psi}(D^{n-k+1}\times D^{k}).$
It yields an oriented bordism from $M$ to $\widetilde{M}$. For simplicity, we will suppress the map $\psi$ from now on in our notation.
Choose a metric $h$ on $T$ that restricts to $g_{M}+\mathrm{d}t^{2}$ on $M\times[0,1]\subset T$. The metric $h$ stays fixed for the remainder of the proof. Choose $R\in(0,\frac{2}{\sqrt{10}}]$ smaller than the normal injectivity radius of $h$ along the core of the surgery
$\mathcal{S}\times[0,1]\;\bigcup\;D^{n-k+1}\times\{0\}\subset T.$
In particular, we assume that $T$ contains the tubular $R$-neighborhood of its core.
Let us choose $\alpha\in(0,\tfrac{1}{3}]$ and let $\varrho_{1}\in(0,\min(R^{2}/4,\alpha/2))$ be the first constant obtained in Theorem 6.1 for $\mathcal{S}\subset(M,g_{M})$. For $0<\varrho\leq\varrho_{1}$, let $B_{\varrho}(D^{n-k+1}\times 0)\subset T$ be the $\varrho$-tubular neighborhood of $D^{n-k+1}\times 0$ in $T$ and let $S_{\varrho}(D^{n-k+1}\times 0)\subset B_{\varrho}(D^{n-k+1}\times 0)$ be the normal sphere bundle of radius $\varrho$ around $D^{n-k+1}\times 0$. Let $h_{S_{\varrho}}$ be the metric on $S_{\varrho}(D^{n-k+1}\times 0)$ induced by $h$.
By decreasing $\alpha$ and hence $\varrho_{1}$, we may assume that for every $0<\varrho\leq\varrho_{1}$, the metric $h_{S_{\varrho}}$ has positive scalar curvature and $S_{\varrho}(D^{n-k+1}\times 0)\subset B_{\varrho}(D^{n-k+1}\times 0)$ has positive mean curvature. For $0<\varrho\leq\varrho_{1}$, let
(7.1) $\Psi_{\varrho_{1},\varrho}\colon S_{\varrho^{4}_{1}}(D^{n-k+1}\times 0)\to S_{\varrho^{4}}(D^{n-k+1}\times 0)$
be the diffeomorphism induced by radial dilatation in the normal bundle of $\mathcal{S}$. By further decreasing $\alpha$, we can assume that the smooth path $(0,\varrho_{1}]\ni\varrho\mapsto\Psi_{\varrho_{1},\varrho}^{*}(h_{\varrho^{4}})$ satisfies
$\partial_{\varrho}\Psi_{\varrho_{1},\varrho}^{*}(h_{\varrho^{4}})\geq-\Psi_{\varrho_{1},\varrho}^{*}(h_{\varrho^{4}}).$
This implies that the smooth path
(7.2) $\varrho\mapsto e^{\varrho}\cdot\Psi_{\varrho_{1},\varrho}^{*}(h_{\varrho^{4}})$
is strictly monotonically increasing.
Let $\sigma\in\mathbb{R}$ and let $(\Omega,g_{\Omega})$ be a fill-in of $(M,g_{M})$ with $\mathrm{scal}(g_{\Omega})\geq\sigma$ and $H_{g_{\Omega}}\geq 0$ on $\partial\Omega$. By a small deformation of $g_{\Omega}$ near $\partial\Omega\subset\Omega$, leaving the induced metric on $\partial\Omega$ unchanged and replacing $\sigma$ with $\sigma-\tfrac{1}{2}$, we can assume that $H_{g_{\Omega}}>0$ along $\partial\Omega$, see [BH23, Proposition 3.8]. By compactness of $\partial\Omega$, there exists a constant $\kappa=\kappa(\Omega,g_{\Omega})>0$ such that $H_{g_{\Omega}}\geq\kappa$.
We choose $\varrho_{0}=\varrho_{0}(\kappa)$ as the second constant from Theorem 6.1. By further decreasing $\varrho_{0}$ we may assume that $(1-\alpha)\kappa>\varrho_{0}^{1/4}$. By choosing $\varrho_{0}$ even smaller, we may assume that
(7.3) $\int_{\partial\Omega}H_{g_{\Omega}}\mathrm{dvol}_{g_{\partial\Omega}}\leq 2\int_{\partial\Omega\setminus B_{4\sqrt{\varrho_{0}}}(\mathcal{S})}H_{g_{\Omega}}\mathrm{dvol}_{g_{\partial\Omega}}.$
Note that $\varrho_{1}$ does not depend on the fill-in $(\Omega,g_{\Omega})$, but $\varrho_{0}$ does.
For the given $\kappa$, consider the metric $g_{\kappa}$ on $M\times[0,1]$ defined in Theorem 6.1,
$g_{\kappa}=u^{\frac{4}{\,n-2\,}}(g_{M}+dt^{2}),\qquad u(x,t)=u_{\kappa}(x,t)=1+t\widetilde{\kappa}\eta\left(\alpha^{-1}t-1\right).$
Recall that $H_{g_{\kappa}}=-\kappa$ along $M\times\{0\}$ with respect to the exterior normal. By Theorem 6.1 and our choice of $\varrho_{0}$, we have
$H(\Sigma_{\varrho_{0}}\cap(M\times[0,1]),g_{\kappa})>(1-\alpha)\kappa-\varrho_{0}^{1/4}>0.$
Since $g_{\kappa}$ restricted to $M\times[\alpha,1]$ equals $g_{M}+\mathrm{d}t^{2}$, the union of $g_{\kappa}$ and $h|_{T\setminus(M\times[0,1])}$ defines a smooth metric $g_{T}$ on $T$. Here, $h$ is the fixed metric on $T$ chosen before. Furthermore, as $g_{\kappa}\stackrel{{\scriptstyle\kappa\to 0}}{{\longrightarrow}}g_{M}+\mathrm{d}t^{2}$ in the $C^{\infty}$-topology, $\mathrm{scal}(T,g_{T})$ is bounded from below by a constant $C(M,g_{M},h)$, which is independent of $\kappa$.
For $0<\varrho\leq\varrho_{1}$ we now define the hypersurface $\Sigma_{\varrho}\subset T$ as follows: The restriction $\Sigma_{\varrho}\cap(M\times[0,1])$ is given by the manifold $\Sigma_{\varrho}\subset M\times[0,1]$ constructed in the beginning of Section 6. Since $a_{\varrho}<2\varrho\leq\alpha$, we see that $\Sigma_{\varrho}\cap\left(M\times[\alpha,1]\right)=S_{\varrho^{4}}(\mathcal{S})\times[\alpha,1]$, where $S_{\varrho^{4}}(\mathcal{S})\subset M$ is the normal sphere bundle of radius $\varrho^{4}$ with respect to the metric $g_{M}$. We extend this by the normal sphere bundle $S_{\varrho^{4}}(D^{n-k+1}\times 0)$ around the upper part $D^{n-k+1}\times\{0\}\subset T$. Let $\Sigma_{\varrho}$ be the resulting smooth manifold and denote the induced Riemannian metric by $g_{\varrho,\kappa}$, see Fig. 11. Since $g_{T}$ is given by the union of $g_{\kappa}$ and $h$, our choice of $\varrho_{1}$ and Theorem 6.1 ensures that the scalar curvature of $(\Sigma_{\varrho},g_{\varrho,\kappa})$ is bounded below by
(7.4) $\mathrm{scal}(M,g_{M})-\varrho_{1}^{1/4}$
for all $0<\varrho\leq\varrho_{1}$ and that the mean curvature of $\Sigma_{\varrho_{0}}$ is positive.
@@
By a smooth deformation of $g_{\Omega}$ we can assure that for some $\delta>0$, the restriction of $g_{\Omega}$ to the tubular neighborhood $B_{2\sqrt{\varrho_{0}}}(\mathcal{S})\times(-\delta,0]\subset\Omega$ of $B_{2\sqrt{\varrho_{0}}}(\mathcal{S})\subset\partial\Omega$ can be smoothly extended by the metric $g_{\kappa}$ onto $B_{2\sqrt{\varrho_{0}}}(\mathcal{S})\times(-\delta,1]$. This can be achieved without altering the induced metric $g_{\varrho,\kappa}$ on $\Sigma_{\varrho}$, $0<\varrho\leq\varrho_{1}$ or the mean curvature of $g_{\Omega}$ on $M\setminus B_{4\sqrt{\varrho_{0}}}(\mathcal{S})$, while preserving $H_{g_{\Omega}}\geq\kappa$. This deformation is described in detail in Section 2, below (2.3), where $U_{i}\subset V_{i}\subset Z_{i}$ correspond to $B_{2\sqrt{\varrho_{0}}}(\mathcal{S})\subset B_{3\sqrt{\varrho_{0}}}(\mathcal{S})\subset B_{4\sqrt{\varrho_{0}}}(\mathcal{S})$.
The manifold $\Sigma_{\varrho_{0}}$ thus bounds an oriented manifold of the form $\Omega_{0}=\Omega\cup T_{0}$, which is diffeomorphic to $\Omega\cup T$ and $g_{T}$ restricts to a smooth metric $g_{\Omega_{0}}$ on $\Omega_{0}$. Hence, we obtain a fill-in $(\Omega_{0},g_{\Omega_{0}})$ of $\Sigma_{\varrho_{0}}$. Since on $M\setminus B_{4\sqrt{\varrho_{0}}}$ the metrics $g_{M}$ and $g_{\varrho_{0},\kappa}$ agree, since $\Sigma_{\varrho_{0}}\subset(\Omega_{0},g_{\Omega_{0}})$ has positive mean curvature and using (7.3), we have
(7.5) $\int_{\partial\Omega}H_{g_{\Omega}}\mathrm{dvol}_{g_{\partial\Omega}}\leq 2\int_{\partial\Omega_{0}}H_{g_{\Omega_{0}}}\mathrm{dvol}_{g_{\partial\Omega_{0}}}.$
In the next step, the right hand side will be bounded above by an application of Proposition 4.2. Let us abbreviate $\Sigma\coloneqq\Sigma_{\varrho_{1}}$ and consider the metric
$g_{\Sigma}\coloneqq e^{\varrho_{1}}\cdot 2^{4/(n-2)}\iota^{*}\big((g_{M}+\mathrm{d}t^{2})\cup h\big),$
where $\iota\colon\Sigma\hookrightarrow T$ denotes the inclusion. Note, that $g_{\Sigma}$ only depends on $(M,g_{M})$ and $\psi$ and that $\Sigma\approx\widetilde{M}$.
We claim that $g_{\varrho_{0},\kappa}\in\mathscr{R}_{g_{\Sigma},s}(\Sigma)$ (see Proposition 4.2) for
$s=-\max(|\mathrm{scal}(M,g_{M})|+\varrho^{1/4})\geq-\max(|\mathrm{scal}(M,g_{M})|-1.$
There exists a smooth path of diffeomorphisms $\Phi_{\varrho}:\Sigma_{\varrho_{1}}\to\Sigma_{\varrho}$, $\varrho_{0}\leq\varrho\leq\varrho_{1}$, such that $\Phi_{\varrho_{1}}=\operatorname{\mathrm{id}}$, $\Phi_{\varrho}|_{M\setminus B_{2\sqrt{\varrho_{1}}}}=\operatorname{\mathrm{id}}$ and such that the path $\varrho\mapsto e^{\varrho}\Phi_{\varrho}^{*}g_{\varrho,\kappa}$ is strictly increasing. On $\Sigma_{\varrho_{1}}\cap(M\times[0,1])$ these diffeomorphisms are constructed in Proposition 6.6. Since the restrictions of these diffeomorphisms to $\Sigma_{\varrho_{1}}\cap(M\times[\tfrac{3}{4},1])$ dilate the normal sphere bundle around $\mathcal{S}\subset M$ from radius $\varrho_{1}^{4}$ to $\varrho^{4}$ (this follows form the choice of $\beta$ in the proof of Proposition 6.6), we can take the union with the dilatations $\Psi_{\varrho_{1},\varrho}$ from (7.1) to define $\Phi_{\varrho_{1},\varrho}$ on $\Sigma_{\varrho}$. By construction, we get $\Phi_{\varrho_{1}}=\operatorname{\mathrm{id}}$, and the path $\varrho\mapsto e^{\varrho}\Phi_{\varrho}^{*}g_{\varrho,\kappa}$ is strictly increasing by Proposition 6.6 and the choice of $\alpha$, see (7.2). This path ends at the metric $e^{\varrho_{1}}g_{\varrho_{1},\kappa}$ on $\Sigma$. Finally, we
  • rescale the conformal factor in $e^{\varrho_{1}}\cdot u^{\tfrac{4}{n-2}}(g_{M}+dt^{2})$, which induces the metric $e^{\varrho_{1}}g_{\varrho_{1},\kappa}$ on $\Sigma\cap(M\times[0,1])$, along the path in Proposition 6.7,
  • rescale the conformal factor in $e^{\varrho_{1}}\cdot h$, which induces the metric $e^{\varrho_{1}}g_{\varrho_{1},\kappa}$ on $\Sigma\setminus(M\times[0,1])$, along the path $s\mapsto e^{\varrho_{1}}(1+s)^{\tfrac{4}{n-2}}$.
This results in a strictly increasing smooth path of metrics on $\Sigma$ starting at $e^{\varrho_{1}}g_{\varrho_{1},\kappa}$ and ending at $g_{\Sigma}$, while maintaining the uniform lower curvature bound $s=-\max(|\mathrm{scal}(M,g_{M})|+\varrho^{1/4})$. This proves our claim that $g_{\varrho_{0},\kappa}\in\mathscr{R}_{g_{\Sigma},s}(\Sigma)$. Let $g_{\Omega_{1}}$ be the smooth Riemannian metric on $\Omega_{1}=\Omega_{0}\cup_{\widetilde{M}}(\widetilde{M}\times[0,1])$ constructed in Proposition 4.2. We have
$\mathrm{scal}_{g_{\Omega_{1}}}\geq\min\{\sigma,s,C(M,g_{M},h)\}-1.$
Put $T_{1}=T_{0}\cup(\widetilde{M}\times[0,1])$ and $C:=\min\{0,s,C(M,g_{M},h)\}$. We obtain
$\int_{\partial\Omega_{0}}H_{g_{\Omega_{0}}}\mathrm{dvol}_{g_{\partial\Omega_{0}}}<\int_{\partial\Omega_{1}}H_{g_{\Omega_{1}}}\mathrm{dvol}_{g_{\partial\Omega_{1}}}.$
Hence the claim follows from (7.5). ∎
Proof Theorem (B).
Let $(M,g_{M})$ and $\sigma\in\mathbb{R}$ be as in Theorem (B). Put $n:=\dim M$. As before, we can assume $n\geq 5$. Since $M$ carries only finitely many spin structures, we can work with a fixed spin structure on $M$.
Since each fill-in of $M$ induces a fill-in of the spin manifold $M\sqcup\overline{M}$, where $\overline{M}$ carries the reverse spin structure, we can furthermore assume that the spin manifold $M$ is spin null-bordant. Let $B$ be a compact spin manifold with $\partial B=M$.
Removing a small open disc from the interior of $B$, we obtain a compact spin bordism $W$ from $M$ to $S^{n}$. Annihilating $\pi_{0}(W)$, $\pi_{1}(W)$ and $\pi_{2}(W)$ by spin surgeries in the interior of $S$ (using $n\geq 5$ and that $W$ is spin), we may assume that the inclusion $S^{n}\hookrightarrow\partial W\subset W$ is $2$-connected. Since $n\geq 5$, this implies that $(W,M)$ has a relative handle decomposition with handles of codimension at least $3$. Hence, the sphere $S^{n}$ is obtained from $M$ by oriented surgeries of codimension at least $3$.
[…]
Remark 7.2.
The construction in Proposition 7.1 allows to impose quantitative restrictions onto the classical surgery step of Gromov–Lawson [GL80]. More specifically, let $\varepsilon>0$, let $\sigma\in\mathbb{R}$, and let $(M,g_{M})$ be a Riemannian manifold satisfying $\mathrm{scal}_{g}>\sigma$. Let $\widetilde{M}$ be obtained from $M$ by surgery along an embedding $\psi\colon S^{n-k}\times D^{k}\hookrightarrow M$, $k\geq 3$. Let $\breve{M}:=M\setminus\mathrm{im}(\psi)$. Choosing $\alpha$ in the proof of Proposition 7.1 small enough, working with a metric $h$ on $T$ of sufficiently small volume and diameter, and working with $M\times[0,\alpha]$ instead of $M\times[0,1]$, we obtain a Riemannian metric $g_{\widetilde{M}}$ on $\widetilde{M}$ with the following properties (compare Eq. 7.4):
  • $\mathrm{scal}_{g_{\widetilde{M}}}>\sigma$,
  • the restriction of $g_{\widetilde{M}}$ to $\breve{M}$ is equal to $g_{M}$,
  • the volume and diameter of the restriction of $g_{\widetilde{M}}$ to $\widetilde{M}\setminus\breve{M}$ are bounded above by $\varepsilon$.